# Decentralized Parametric Stability Certificates for Grid-Forming Converter Control

### v5

@@ 3.2 Grid-Forming Converter Dynamics
Fig. 3 shows the basic implementation of a GFM three-phase VSC connected to the power network in Fig. 2. The linearized small-signal dynamics (in the global per unit system) of the $i$-th VSC ($i\in\{1,...,n\}$) can be represented by the $2\times 2$ transfer matrix $D_{i}(s)$, mapping from the active and reactive power injections $\Delta p_{i}(s)$ and $\Delta q_{i}(s)$ to the imposed frequency and voltage magnitude $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$, i.e.,
@@
The most prevalent GFM control strategy is filtered droop control [21]. Assuming time-scale separation, we can neglect the internal dynamics of the VSC (see Appendix 8 for a detailed derivation) and consider $D_{i}(s)$ as in the following definition. Of course, all our developments will be evaluated on the full VSC model in Section 5.
@@
The small-signal converter dynamics 20 used to derive the stability certificates in Section 4 are given by the droop controller
@@
Finally, if we extend 20 to include the dynamics of all $n$ converters, we obtain the following $2n\times 2n$ transfer matrix
$-\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\omega(s)\\ \Delta|v|(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\omega_{1}(s)\\ \Delta{|v|}_{1}(s)\\ \vdots\\ \Delta\omega_{n}(s)\\ \Delta{|v|}_{n}(s)\end{bmatrix}}}=\underset{\eqqcolon\,D(s)}{\underbrace{\begin{bmatrix}D_{1}(s)&0_{2\times 2}&\dots&0_{2\times 2}\\ 0_{2\times 2}&D_{2}(s)&\dots&0_{2\times 2}\\ \vdots&\vdots&\ddots&\vdots\\ 0_{2\times 2}&0_{2\times 2}&\dots&D_{n}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}.$ (22)
$\mathcal{D}_{0}(s)$$\mathcal{N}_{0}(s)$--$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|v|\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|v|_{\mathrm{d}}\end{bmatrix}$device dynamicsnetwork dynamics--$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
@@ 3.3 Dynamic Power System Model
The closed-loop power system dynamics are modeled as the feedback interconnection of the converter device dynamics in 22 and the network dynamics in 11 as illustrated in Fig. 4. We focus on the stability of the bus frequency and voltage magnitudes, given by $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$ for $i\in\{1,...,n\}$ and consider these quantities as interconnection signals between the device and network dynamics in Fig. 4. Accordingly, the interconnected subsystems are defined as:
@@
$c_{i,\rho}$ Definition $c_{i,\rho}$ Definition
$c_{1,\rho}$ $({1+\rho^{2}})/({5+2\rho^{2}})$ $c_{8,\rho}$ $2\rho\omega_{0}(15|v|_{\mathrm{max}}^{2}+4)(10\rho^{2}|v|_{\mathrm{max}}^{2}+25|v|_{\mathrm{max}}^{2}+4)$
$c_{2,\rho}$ ${(1+\rho^{2})^{2}}/({6\rho})$ $c_{9,\rho}$ $10\omega_{0}\rho^{4}|v|_{\mathrm{max}}^{2}+95\omega_{0}\rho^{2}|v|_{\mathrm{max}}^{2}+16\omega_{0}\rho^{2}+25\omega_{0}|v|_{\mathrm{max}}^{2}$
$c_{3,\rho}$ $2\rho(1+\rho^{2})$ $c_{10,\rho}$ $5\omega_{0}(1+\rho^{2})(20\rho^{3}|v|_{\mathrm{max}}^{2}+80\rho|v|_{\mathrm{max}}^{2}+16\rho)$
$c_{4,\rho}$ $((2\rho^{2}-5)+\sqrt{(5-2\rho^{2})^{2}+16\rho^{2}})/4\rho$ $c_{11,\rho}$ $50\rho\omega_{0}(1+\rho^{2})^{2}$
$c_{5,\rho}$ ${(1+\rho^{2})}/({10\rho^{2}|v|_{\mathrm{max}}^{2}+25|v|_{\mathrm{max}}^{2}+4)}$ $c_{12,\rho}$ ${(\rho^{2}+1)^{2}}/({6\rho|v|_{\mathrm{max}}^{2}})$
$c_{6,\rho}$ ${5(1+\rho^{2})^{2}}/({30\rho|v|_{\mathrm{max}}^{2}+8\rho)}$ $c_{13,\rho}$ $-2\rho|v|_{\mathrm{max}}^{4}$
$c_{7,\rho}$ ${5(1+\rho^{2})}/{4}$ $c_{14,\rho}$ $|v|_{\mathrm{max}}^{2}(2\rho^{2}-5)$
Table 1: Definition of the quantities $c_{l,\rho}$ for $l=1,...,14$ in the conditions 24 and 25 depending on the parameter $\rho$.
@@
Figure 5: Graphical illustration of the stability conditions in 24, where $\alpha_{\mathrm{p},i}=\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{p},i}=\tau_{\mathrm{p},i}\omega_{0}$,.
@@
Figure 6: Graphical illustration of the stability conditions in 25, where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$.
4 Decentralized Stability Conditions
By applying dynamic loop-shifting techniques and passivity theory (see Section 4.2 for details), we derive parametric decentralized stability conditions that serve as local tuning rules for the dynamic droop control 21 of each VSC $i\in\{1,\dots,n\}$, ensuring the internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig. 4. The stability conditions for each VSC $i$ depend on
Crucially, the conditions are local and entirely independent of the control parameters of other VSCs, making them directly applicable for scalable and decentralized stability assessment, as well as for local controller tuning and grid-code design.
The decentralized stability conditions for each VSC $i$ can be categorized into decoupled constraints on the active power-frequency droop control and the reactive power-voltage droop control. For each control scheme, these conditions can be visualized in either a two-dimensional (for fixed $\rho$) or a three-dimensional (for variable $\rho$) coordinate system, as shown in Figures 5 and 6. The coordinate axes represent scaled versions of the local droop gains and time constants (in 2D) and the global network parameter $\rho$ (in 3D), with typical value ranges. The resulting feasible parameter sets are indicated by green dots. As illustrated in Fig. 5, for small values of $\rho$, closed-loop stability is ensured when the droop gain $d_{\mathrm{p},i}$ and/or the self-susceptance $\sum_{j\neq i}^{n}b_{ij}$ are sufficiently small, irrespective of the time constant $\tau_{\mathrm{p},i}$. Notably, this suggests potential instability issues in future transmission systems with a high density of devices interconnected by short lines, which correspond to large values of $\sum_{j\neq i}^{n}b_{ij}$. Further, interpreting the filter time constant $\tau_{\mathrm{p},i}$ as virtual inertia reveals that increasing virtual inertia (i.e., larger $\tau_{\mathrm{p},i}$) does not necessarily enhance system stability. These findings align with the small-signal stability conditions derived for active power droop control in conjunction with SISO frequency dynamics in [8].
Finally, from the 3D plots in Figures 5 and 6, it becomes apparent how an increasing $\rho$ allows for larger local droop gains of both the active and reactive power droop control.
In what follows, we provide an algebraic parametrization of the conditions in Figures 5 and 6 (a detailed derivation is provided in Section 4.2). Namely, for the active power-frequency droop control, we require
@@ 4 Decentralized Stability Conditions
where $\alpha_{\mathrm{p},i}=\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{p},i}=\tau_{\mathrm{p},i}\omega_{0}$, and $c_{l\rho}$ for $l=\{1,2,3,4\}$ are quantities depending on $\rho$ as defined in Table 1. For the reactive power-voltage droop control, we require
$\begin{split}\alpha_{\mathrm{q},i}&<c_{5,\rho}(5+10\rho\tilde{\tau}_{\mathrm{q},i})\end{split}$ (25a)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{6,\rho}(2\rho+\tilde{\tau}_{\mathrm{q},i})\end{split}$ (25b)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{7,\rho}\end{split}$ (25c)
$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{8\rho}-\alpha_{\mathrm{q},i}(5\tilde{\tau}_{\mathrm{q},i}(1+\rho^{2})c_{9,\rho}+c_{10,\rho})\\ &\quad\quad\quad+c_{11,\rho}(\tilde{\tau}_{\mathrm{q},i}^{2}(1+\rho^{2})+2\rho\tilde{\tau}_{\mathrm{q},i}+1)\end{split}$ (25d)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{12,\rho}\tilde{\tau}_{\mathrm{q},i}\end{split}$ (25e)
$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{13,\rho}+\alpha_{\mathrm{q},i}\tilde{\tau}_{\mathrm{q},i}c_{14,\rho}+2\tilde{\tau}_{\mathrm{q},i}^{2}\rho\end{split}$ (25f)
@@
The conditions in 24 and 25 are designed for the scenario where all $n$ converter nodes are populated, ensuring system-level stability during disconnection and re-connection of VSCs. Specifically, removing any VSC decreases the self susceptance $\sum_{j\neq i}^{n}b_{ij}$ for each connected VSC $i$ after Kron reduction and reindexing, thereby guaranteeing that 24 and 25 remain locally satisfied.
[…]
Figure 7: Structural overview of the proof of Theorem 2.
@@ 4.1 Main Result
The conditions 24 and 25 are derived in Section 4.2, which proves our main result stated in the theorem below:
@@
Consider the device and network dynamics $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 23, where the network model is given by 17. Let the conditions in 24 and 25 hold for each VSC $i\in\{1,...,n\}$. Then, the closed-loop system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig. 4 is internally feedback stable, i.e., $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
@@
For the quasi-stationary network dynamics $\mathcal{N}_{0}(s)$ with the network model in 18, where $s=0$ and $|v|_{0,i}\neq|v|_{0,j}$, the stability result of Theorem 2 holds under the relaxed algebraic conditions [^1] These conditions structurally align with the conditions derived in [9].
$\begin{split}\alpha_{\mathrm{q},i}<\tfrac{5(1+\rho^{2})}{4}=c_{7,\rho}\quad\text{and}\quad 0<\tau_{\mathrm{p},i}\end{split}$ (26)
for each VSC $i\in\{1,...,n\}$. Further, for the zero-power-flow network dynamics $\mathcal{N}_{0}(s)$ with the network model in 19, where $s=0$ and $|v|_{0,i}=|v|_{0,j}$, internal feedback stability is guaranteed without imposing any additional VSC conditions.
The proof of Corollary 1 is provided in Appendix 9 and follows a similar reasoning as the proof of Theorem 2 in Section 4.2, but applies the simplified network models in 18 and 19, respectively. It becomes apparent how the conditions in 24 and 25 are a subset of the conditions in 26.

### v6

@@ 3.2 Grid-Forming Converter Dynamics
Fig. 3 shows the implementation of a GFM three-phase VSC [^1] A simple three-phase VSC implementation is used here for illustration. However, our results are not limited to this specific converter architecture and can be extended to other filter configurations and inner-loop structures, provided the time-scale separation discussed in Remark 4 holds. connected to the power network in Fig. 2. The linearized small-signal dynamics (in the global per unit system) of the $i$-th VSC ($i\in\{1,...,n\}$) can be represented by the $2\times 2$ transfer matrix $D_{i}(s)$, mapping from active and reactive power injections $\Delta p_{i}(s)$ and $\Delta q_{i}(s)$ to the imposed frequency and voltage magnitude $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$, i.e.,
@@
The most prevalent GFM control strategy is filtered droop control [23]. Assuming time-scale separation as elaborated in Remark 4, we can neglect the internal dynamics of the VSC (see Appendix 8 for a derivation) and consider $D_{i}(s)$ as in the following definition. Of course, all our developments will be evaluated on the full VSC model in Section 5.
@@
The small-signal converter dynamics 21 used to derive the stability certificates in Section 4 are given by the droop controller
@@
Finally, if we extend 21 to include the dynamics of all $n$ converters, we obtain the following $2n\times 2n$ transfer matrix
[…]
Remark 4.
From a power electronics perspective, three distinct dynamic time scales are typically considered (see Fig. 5):
  • (i)
    fast inner VSC control loops,
  • (ii)
    intermediate network dynamics, and
  • (iii)
    slow outer VSC control loops.
These layers must be sufficiently separated in this order; otherwise, standard assumptions for control design and model reduction break down. Consistent with [9, 12], in this work, we assume that inner-loop and network dynamics are well separated, while a possible overlap between network and outer-loop dynamics is treated as a separate problem.
$\mathcal{D}_{0}(s)$$\mathcal{N}_{0}(s)$-$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|v|\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|v|_{\mathrm{d}}\end{bmatrix}$device dynamicsnetwork dynamics-$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
@@
1msphysicsconrolinner convertercontrol (current+ voltage loop)outer convertercontrol (active &reactive power loopsnetworkdynamicstimescales of interest10ms100ms
Figure 5: Time scales of interest for our stability study, adapted from [12].
@@ 3.3 Dynamic Power System Model
The closed-loop power system dynamics are modeled as the feedback interconnection of the converter device dynamics in 23 and the network dynamics in 12 as illustrated in Fig. 4222An extension to heterogeneous device allocations with synchronous generators, grid-forming and -following converters will be part of future work.. We focus on the stability of the bus frequency and voltage magnitudes, given by $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$ for $i\in\{1,...,n\}$ and consider these quantities as interconnection signals between the device and network dynamics in Fig. 4. Accordingly, the interconnected subsystems are defined as:
@@
$\hskip-2.84526ptc_{i,\rho}\hskip-2.84526pt$ Definition $\hskip-2.84526ptc_{i,\rho}\hskip-2.84526pt$ Definition
$\hskip-2.84526ptc_{1,\rho}\hskip-2.84526pt$ $({1+\rho^{2}})/({5+2\rho^{2}})$ $c_{6,\rho}\hskip-2.84526pt$ ${(1+\rho^{2})}/({2\rho^{2}|v|_{\mathrm{max}}^{2}+5|v|_{\mathrm{max}}^{2})}$
$\hskip-2.84526ptc_{2,\rho}\hskip-2.84526pt$ ${(1+\rho^{2})^{2}}/({6\rho})$ $c_{7,\rho}$ ${(\rho^{2}+1)^{2}}/({6\rho|v|_{\mathrm{max}}^{2}})$
$\hskip-2.84526ptc_{3,\rho}\hskip-2.84526pt$ $2\rho(1+\rho^{2})$ $c_{8,\rho}$ $2\rho|v|_{\mathrm{max}}^{4}$
$\hskip-2.84526ptc_{4,\rho}\hskip-2.84526pt$ $((2\rho^{2}-5)+\sqrt{(5-2\rho^{2})^{2}+16\rho^{2}})/4\rho$ $c_{9,\rho}$ $|v|_{\mathrm{max}}^{2}(2\rho^{2}-5)$
$\hskip-2.84526ptc_{5,\rho}\hskip-2.84526pt$ ${5(1+\rho^{2})}/{4}$
Table 1: Definition of the quantities $c_{l,\rho}$ for l=1,,9l=1,...,9 in the conditions 25 and 26 depending on the parameter $\rho$.
4 Decentralized Stability Conditions
In this section, we first introduce the decentralized stability conditions, which provide practical guidelines for local converter tuning and decentralized stability assessment. We then present our main result, which guarantees internal feedback stability of the closed-loop system, assuming the previously stated stability conditions are satisfied. Notably, these conditions naturally emerge from the proof of the main result.
4.1 Decentralized Stability Conditions
By applying dynamic loop-shifting techniques and passivity theory (see Section 4.3 for details), we derive parametric decentralized stability conditions that serve as local tuning rules for the dynamic droop control 22 of each VSC $i\in\{1,\dots,n\}$, ensuring the internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig. 4. The stability conditions for each VSC $i$ depend on
Crucially, the conditions are local and entirely independent of the control parameters of other VSCs, making them directly applicable for scalable and decentralized stability assessment, as well as for local controller tuning and grid-code design. Beyond that, the decentralized stability certificates support heterogeneous device-level controllers by allowing each converter to be designed with independent droop gains and time constants, without requiring uniform parameters across the system, as is the case in, e.g., [9].
The decentralized stability conditions for each VSC $i$ can be categorized into decoupled constraints on the active power-frequency droop control and the reactive power-voltage droop control. For each control scheme, these conditions can be visualized in either a two-dimensional (for fixed $\rho$) or a three-dimensional (for variable $\rho$) coordinate system, as shown in Figures 6 and 7. The coordinate axes represent scaled versions of the local droop gains and time constants (in 2D) and the global network parameter $\rho$ (in 3D), with typical value ranges. The resulting feasible parameter sets are indicated by green dots. As illustrated in Fig. 6, for small values of $\rho$, closed-loop stability is ensured when the droop gain $d_{\mathrm{p},i}$ and/or the self-susceptance $\sum_{j\neq i}^{n}b_{ij}$ are sufficiently small, irrespective of the time constant $\tau_{\mathrm{p},i}$. Notably, this suggests potential instability issues in future transmission systems with a high density of devices interconnected by short lines, which correspond to large values of $\sum_{j\neq i}^{n}b_{ij}$. Further, interpreting the filter time constant $\tau_{\mathrm{p},i}$ as virtual inertia reveals that increasing virtual inertia (i.e., larger $\tau_{\mathrm{p},i}$) does not necessarily enhance system stability. These findings align with the small-signal stability conditions derived for active power droop control in conjunction with SISO frequency dynamics in [9].
Finally, from the 3D plots in Figures 6 and 7, it becomes apparent how an increasing $\rho$ allows for larger local droop gains of both the active and reactive power droop control.
@@ 4.1 Decentralized Stability Conditions
Figure 6: Graphical illustration of the stability conditions in 25, where $\alpha_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tau_{\mathrm{p},i}\omega_{0}$ (in rad).
@@
Figure 7: Graphical illustration of the stability conditions in 26, where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$ (in rad).
In what follows, we provide an algebraic parametrization of the conditions in Figures 6 and 7 (a detailed derivation is provided in Section 4.3). Namely, for the active power-frequency droop control, we require
@@
where $\alpha_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tau_{\mathrm{p},i}\omega_{0}$, and $c_{l\rho}$ for $l\hskip-1.42262pt=\hskip-1.42262pt\{1,2,3,4\}$ are quantities depending on $\rho$ as defined in Table 1. For the reactive power-voltage droop control, we require
$\begin{split}\alpha_{\mathrm{q},i}&<c_{5,\rho}\end{split}$ (26a)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{6,\rho}2\rho\tilde{\tau}_{\mathrm{q},i}\end{split}$ (26b)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{7,\rho}\tilde{\tau}_{\mathrm{q},i}\end{split}$ (26c)
$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{8,\rho}+\alpha_{\mathrm{q},i}\tilde{\tau}_{\mathrm{q},i}c_{9,\rho}+2\tilde{\tau}_{\mathrm{q},i}^{2}\rho\end{split}$ (26d)
@@
The conditions in 25 and 26 are designed for the scenario where all $n$ converter nodes are populated, ensuring system-level stability during disconnection and re-connection of VSCs. Specifically, removing any VSC decreases the self susceptance $\sum_{j\neq i}^{n}b_{ij}$ for each connected VSC $i$ after Kron reduction and reindexing, thereby guaranteeing that 25 and 26 remain locally satisfied.
@@ 4.2 Main Result
The conditions 25 and 26 are derived in Section 4.3, which proves our main result stated in the theorem below:
@@
Consider the device and network dynamics $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 24, where the network model is given by 18. Let the conditions in 25 and 26 hold for each VSC $i\in\{1,...,n\}$. Then, the closed-loop system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig. 4 is internally feedback stable, i.e., $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
@@
For the quasi-stationary network dynamics $\mathcal{N}_{0}(s)$ with the network model in 19, where $s=0$ and $|v|_{0,i}\neq|v|_{0,j}$, the stability result of Theorem 2 holds under the relaxed algebraic conditions
$\begin{split}\alpha_{\mathrm{q},i}<\tfrac{5(1+\rho^{2})}{4}=c_{5,\rho}\quad\text{and}\quad 0<\tau_{\mathrm{p},i}\end{split}$ (27)
for each VSC $i\in\{1,...,n\}$. Further, for the zero-power-flow network dynamics $\mathcal{N}_{0}(s)$ with the network model in 20, where $s=0$ and $|v|_{0,i}=|v|_{0,j}$, internal feedback stability is guaranteed without imposing any additional VSC conditions.
The proof of Corollary 1 is provided in Appendix 9 and follows a similar reasoning as the proof of Theorem 2 in Section 4.3. It becomes apparent how the conditions in 25 and 26 are a subset of the conditions in 27. The latter structurally align with the conditions derived in [10].